if a+3d/a+9d=a+d/a+5d=k,then k=(a,d>0) solution

Grade:11

1 Answers

Vikas TU
7 years ago
Let a+3d/a+9d=a+d/a+5d=k
then,
          a2+5ad+3ad+15d2= a2+9ad+ad+9d2=k
          8ad+15d2=10ad+9d2=k
          6d2=2ad=k
          2a=6d=k
          a=3d=k
Hnece k = 3d is the solution.